OpenStudy (anonymous):

For the function f(x) = 3x+2, find x such that f(x) = 14.

OpenStudy (anonymous):

Alright, let's do this! Love your name btw

OpenStudy (anonymous):

well so 14 = 3x + 2

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Ok so let's evaluate f(x) at 14 This is f(14) So we plug in 14 for the x in the function and evaluate! Like so... f(14) = 3(14)+2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Do you know what to do now?

OpenStudy (anonymous):

i get what you are saying @Cosmichaotic and ya i do let me do the work and ill ask if the answer is right

OpenStudy (haseeb96):

f(x)=3x+2 and f(x)=14 u have found one thing common in both the equation that is the left hand term is same so put the 14 there then it become 14=3x+2 3x+2-14=0 3x-12=0 3x=12 x=12/3 x=4 is the correct answer

OpenStudy (anonymous):

@Haseeb96 is right ;)

OpenStudy (anonymous):

Oh, where f(x) = 14, not for x = 14, my bad.

OpenStudy (anonymous):

i got 44

OpenStudy (anonymous):

no @Shadowgirl15 forget that...

OpenStudy (haseeb96):

No @PFEH.1999 is correct my brother @Cosmichaotic

OpenStudy (anonymous):

so the answer is 4?

OpenStudy (haseeb96):

@Shadowgirl15 please have serve of my explanation

OpenStudy (anonymous):

We are actually setting f(x) to 14 here, and not x to 14 So we need to say f(x) = 14 f(x) = 3x+2 So 14 = 3x + 2 and then solve for x.

OpenStudy (haseeb96):

yes

OpenStudy (anonymous):

Thanks!!!

OpenStudy (anonymous):

holy crap lloks like i came here fo no reason lol

OpenStudy (haseeb96):

most welcome my dear